OpenStudy (anonymous):

The rectangles below are similar. The length of the larger rectangle is 3x + 1 and the width is 8. The length of the smaller rectangle is 5 and the width is 4. What is the length of the larger rectangle?

OpenStudy (dumbcow):

similar rectangles are proportional. Meaning the ratio of length/width is same \[\frac{3x+1}{8} = \frac{5}{4}\] solve for x

OpenStudy (anonymous):

3